1. A homogeneous products duopoly, Paper co and Wow Paper Inc, is considering new market in Brazil and it faces a market demand function given P = 300 – 3Q, where Q=Q1 +Q2. Both firms have a constant marginal cost MC = 100.

e) Since Paper Co has timber production for the paper in Amazon, their MC to produce paper would be lower than Wow Paper Inc. What are the Cournot equilibrium quantities and industry price when one firm (Wow Paper Inc) has a marginal cost of 100 but the other firm (Paper Co) has a marginal cost of 90?

f)Support your answer by showing the game table with associated payoff.
Answer #1

E)p=300-3Q1-3Q2

MR1=300-6Q1-3Q2

MC=100

MR1=MC1

200-3Q2=6Q1

Q1=100/3-0.5Q2{ best response of firm 1 with MC=100( wow paper)

MR2=300-6Q2-3Q1

MC2=90

MR2=MC2

210-3Q1=6Q2

Q2=35-0.5 Q1{ best response of paper co}

Putting q1 into q2,

Q2=35-0.5(100/3-0.5q2)

Q2=35-50/3+0.25q2

Q2=(55/3)*1/0.75=(55/3)*4/3=220/9

Q1=100/3-0.5*220/9=100/3-110/9=190/9

Q=220/9+190/9=410/9

P=300-3*410/9=300-410/3=490/3

F)Date firm 2 790/9 220/9 190/91848 2020.02.12 08:20 Cournot 7548 X1193 Nasty bru esne m2 1760 /11337 1548 2148/1304 R1548